I need help with story problems. Please Help!
There suppose to be easy, but I don't see how.
1. There are 3 children in my family and our total ages up to 34. I am 4 times as old as my brother, and my younger sister is 4 years older than my brother. How old am I?
2. Jamie has 4 more points than Sabrina,. If they have 56 points between them, how many points does each have?
3. The length of a large rectangular room is twice the width. If the perimeter is 36 yards, find the dimensions of the room.
4. One day Tom, Bob, and Jane sat down to figure out who was oldest. Their combined ages came to 54. Tom is four times as old as Bob less 10. Bob is four less than Jane. Who is the oldest, and how old are they?
PLEASE, PLEASE, PLEASE HELP!!!!

Answer by longjonsilver(2297)

1. Let brother's be age x, since the brother is mentioned in both other ages.

3 ages are therefore: x and 4x and x+4

x+4x+x+4 = 34
6x+4 = 34
6x = 30
x = 30/6
x = 5

So brother is 5. Person talking is therefore 20.

2. Let sabrina have p points so that Jamie has p+4.

So, p + p+4 = 56
2p + 4 = 56
2p = 52
p = 52/2
p = 26

Sabrina had 26 points and Jamie had 30.

3. Let width = w. So length = 2w

perimeter is w + w + 2w + 2w = 36
6w = 36
w = 36/6
--> width = 6 yards
--> length = 12 yards

4. you try it.

jon.